OpenStudy (anonymous):

If the line y=20x+1 intersects the curve y=b^x, b>1, and x=4, then b= A)2 B)3 C)4 D)5 E)6 please explain

OpenStudy (anonymous):

u see my predicament

OpenStudy (anonymous):

I break it down: 1) y = 20x+1 line x =4 intersects at???

OpenStudy (anonymous):

i think 81?

OpenStudy (anonymous):

your answer is y =81. Actually, they intersect at the point (4,81) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

now 2) y = b^x and you have 81 = b^4, right?

OpenStudy (anonymous):

im so stupid

OpenStudy (anonymous):

never noticed before thank you

OpenStudy (anonymous):

B

OpenStudy (anonymous):

its fckin B

OpenStudy (anonymous):

wonder how it is c?????

OpenStudy (anonymous):

oh yeah, better now hihihihi

OpenStudy (anonymous):

3^4=81

OpenStudy (anonymous):

derp

OpenStudy (anonymous):

you cheat!! you said it is c, then delete and repost it's B. hihihi whatsoever, B is correct answer

OpenStudy (anonymous):

what